Write an equation of the line that passess through the given points. (2,12), (7,2)
Answer by checkley71(8403)

FIRST FIND THE SLOPE (Y2-Y1)/(X2-X1)
(2-12)/(7-2)
-10/5
-2 IS THE SLOPE (m).
NOW SUBSTITUTE ONE SET OF POINTS IN THE LINE FORMULA & SOLVE FOR THE Y INTERCEPT (b) IN THE FORMULA
Y=mX+b
12=-2*2+b
12=-4+b
b=12+4
b=16 THUS THE FORMULA IS
Y=-2X+16